6. Determine the sum of the first 75 terms of the sequence 4, 1, -2, -5, -8,...
Answer #1

SOLUTION;

the series number is

P= 4,1,-2,-5,-8,........

a= 4 =first number

d= different between two corresponding number

d= 1-4=-2-1=-5-(-2)=........=-3

And fund first 75 number sum so

n=75

We know that, sum of n first number of series is

S_n=\frac{n}{2}[2a+(n-1)d]

S_{75}=\frac{75}{2}[2a+(75-1)d]

S_{75}=\frac{75}{2}[2(4)+(75-1)(-3)]

S_{75}=\frac{75}{2}[8+(74)(-3)]

S_{75}=\frac{75}{2}(-214)

\therefore S_{75}=-8025

The sum of first 75 number is S_{75}=-8025

Answer =-8025